# LLM-Based Multi-Agent Systems for Software Engineering: Vision and the Road Ahead

### v2

@@ 4. Research Agenda
Previous research has laid the groundwork for the exploration of LMA systems in software engineering, yet this domain remains in its nascent stages, with many critical challenges awaiting resolution. In this section, we outline our perspective on these challenges and suggest research questions that could advance this burgeoning field. We envision two phases for the development of LMA systems in software engineering. We discuss each of these phases below and suggest a series of research questions that could form the basis of future research projects.
@@ 4.1.1. Refining Role-Playing Capabilities in Software Engineering
The role-playing capabilities of LLM-based agents are pivotal within an LMA system (Wang et al., 2023d). Given the complex nature of SE tasks, we require specialized agents to focus on specific information, apply specialized knowledge, and think from diverse perspectives to generate meaningful and effective outputs. While LLM-based agents have shown proficiency in simulating well-recognized roles like generic software developers or project managers (Hong et al., 2023; Qian et al., 2024c), they often struggle to accurately represent specialized expert roles that demand extensive domain expertise. For example, roles involving vulnerability detection or security auditing require a nuanced understanding of security protocols, threat modeling, and the latest vulnerabilities, which goes beyond general programming knowledge. Similarly, roles like database optimization specialists or embedded systems engineers demand specialized skills and insights that are challenging for LLMs to emulate accurately. This shortfall arises primarily because they are trained on general text corpora that lack the specialized content necessary to grasp the nuanced expertise required in SE (Hou et al., 2023). This limitation highlights the need for integrating domain-specific expertise into LLMs to better support specialized roles within software engineering. To address this limitation, we propose a structured and actionable three-step approach encompassing the identification, assessment, and enhancement of role-playing abilities, which are:
@@
    Value Addition Modeling: The next crucial step is value addition modeling (Mendes et al., 2018), which evaluates the potential advantages that LLM-based agents could bring to each prioritized role. This process involves constructing detailed, data-driven models to analyze key performance indicators such as efficiency improvements, cost reductions, quality enhancements, and the acceleration of innovation resulting from the integration of agents. Pilot projects can be deployed to gather empirical data on these metrics when LLM-based agents are applied to specific tasks. Important factors to consider include the automation of repetitive tasks, the augmentation of human capabilities, and the enablement of new functionalities that were previously unattainable. It is important to note that the value added by LLM-based agents can differ significantly across different domains; for example, roles in software development may prioritize automation, whereas domains like systems architecture might see more value in LLMs augmenting complex decision-making around resource allocation or performance optimization, where human expertise and contextual understanding remain essential. By quantifying these value propositions, organizations can allocate resources more strategically to roles where LLM-based agents are likely to yield the highest return on investment.
@@ 4.1.2. Advancing Prompts through Agent-oriented Programming Paradigms
In previous sections, we discussed the benefits of well-crafted prompts. However, crafting such prompts is indeed challenging due to the need for a versatile, effective, and robust framework applicable across various scenarios. Natural language, while flexible, often contains ambiguities and inconsistencies that LLMs may misinterpret. This highlights the need for a specialized prompting language designed to augment the cognitive functions of LLM-based agents. Natural language is inherently designed for human communication, relying on shared context and intuition that LLMs lack. LLMs interpret text based on statistical patterns from large datasets, which can lead to different interpretations than those intended for humans (Sun et al., 2024a; Zeng et al., 2022). To bridge this gap, we propose the development of a formalized prompting language that treats LLMs as the primary audience. Such a language aims to minimize ambiguities and ensure clear, deterministic instructions, resulting in more reliable and accurate outputs.
@@
Agent-oriented programming (AOP) (Shoham, 1993) offers a promising foundation for this approach. Similar to how Object-Oriented Programming (OOP) (Wegner, 1990) organize objects, AOP treats agents as fundamental units, focusing on their reasoning, objectives, and interactions. An AOP-based prompting language could enable the precise expression of complex tasks and constraints, allowing LLM-based agents to perform their roles with greater efficiency and accuracy. By clearly defining instructions within the AOP framework, we can reduce ambiguity, mitigate hallucinations, and enhance task execution.
However, creating a new language for prompts presents its own challenges. Such a language must be expressive enough to handle diverse and complex tasks, yet simple enough for users to easily adopt. Balancing functionality and usability will be critical to its success. Additionally, this process may involve tailoring prompts specifically for different LLM models and their versions, as variations in model architectures, training data, and capabilities can affect how they interpret and respond to prompts. What works effectively for one model may not perform as well for another, necessitating careful adjustments. While AOP-based prompting may not be the final solution, it represents an important step toward developing an AI-oriented language with grammar tailored specifically for LLMs. This new approach could further refine communication with LLM-based agents, reducing misinterpretation and significantly enhancing overall performance.
@@ 4.2.1. Human-Agent Collaboration
Optimally distributing tasks between humans and LMA agents to leverage their respective strengths is essential. Humans bring unparalleled creativity, critical thinking, ethical judgment, and domain-specific knowledge (Markauskaite et al., 2022). In contrast, LMA agents excel at rapidly processing large datasets, performing repetitive tasks with high accuracy, and detecting patterns that might elude human observers. Aligning LMA systems with human intentions and expertise is critical, but integrating human involvement throughout the software development lifecycle remains insufficiently explored. Identifying the most effective roles for humans, optimizing feedback mechanisms, and determining the appropriate moments for human intervention present significant challenges.
To facilitate seamless collaboration, designing intuitive user-friendly interfaces and interaction protocols is essential (Wang et al., 2024a). Natural language interfaces and adaptive visualization techniques can make interactions more accessible. These interfaces should efficiently present agent outputs in a digestible format and collect user feedback, while also managing the cognitive load on human collaborators. It is important to note that these interfaces may need to be tailored differently for each human role, as the needs of a project manager, a software developer, and a quality assurance engineer will vary significantly. Given the complexity of information generated during the agents’ workflows (Josifoski et al., 2023), designing such interfaces poses challenges. For instance, presenting modifications suggested by an agent at varying levels of abstraction ensures that each stakeholder can engage with the information at the right depth. A project manager might focus on the broader implications, such as the high-level impact on project timelines or deliverables, whereas a developer or architect might drill down into specific implementation details. Role-specific interfaces will be key to ensuring each stakeholder can effectively collaborate with the agents and extract the necessary information in a manner suited to their specific responsibilities.
@@ 4.2.2. Evaluating the LMA systems
Recently, numerous complex benchmarks have been proposed to challenge the capabilities of LLMs in critical aspects of software development, such as code generation (Zhuo et al., 2024; Jain et al., 2024; Liu et al., 2024c). While these benchmarks have advanced the field by providing measurable metrics for individual tasks, their focus on isolated problem-solving reveals limitations as software engineering projects become more complex.
Software engineering is inherently collaborative, with key activities like joint requirements gathering, code integration, and peer reviews playing essential roles in the process. Current benchmarks often overlook these aspects, failing to assess how well LLMs perform in tasks that require cooperation and collective decision-making. There is a growing need for benchmarks that evaluate the cooperative abilities of LLMs in multi-agent settings, particularly for software engineering tasks. These benchmarks should simulate real-world collaborative scenarios where LLM agents work together to achieve common development goals.
@@
First, as software projects grow in size and complexity, breaking down high-level requirements into manageable sub-tasks becomes more difficult. It is not just about handling more tasks but also managing the intricate interdependencies between them. A hierarchical task decomposition approach can help, where higher-level agents oversee broader objectives and delegate specific tasks to lower-level agents. This structure streamlines planning and makes global task allocation more efficient.
@@ 4.2.4. Leveraging Industry Principles
In the context of software development, predicting the optimal configuration for LMA systems at the outset is unrealistic due to the inherent complexity and variability of tasks (Leffingwell and Widrig, 2000). The dynamic nature of software requirements and the unpredictable challenges that arise during development necessitate systems that can adapt on the fly (Liu et al., 2023). Therefore, LMA systems must be capable of dynamically adjusting their scale, strategies, and structures throughout the development process.
To minimize redundant work, LMA systems should continuously evaluate existing solutions (Wang et al., 2024c), identifying reusable elements for new requirements. By learning from each development cycle, the system can recognize patterns of efficiency and inefficiency, enabling it to make informed decisions when handling similar tasks in the future or adapting existing solutions to new requirements.
@@
To address these challenges, robust and fine-grained access control mechanisms must be implemented across organizational boundaries. It is essential to prevent unauthorized access while still accommodating the varied data access needs of the system. Traditional models like Role-Based Access Control (RBAC) (Sandhu, 1998) and Attribute-Based Access Control (ABAC) (Hu et al., 2015) may need to be extended to handle the dynamic nature of multi-agent systems effectively. Establishing protocols that allow agents to share insights derived from sensitive data, without exposing the data itself, is critical. Advanced privacy-preserving techniques like Differential Privacy (Dwork, 2006), Secure Multi-Party Computation (SMPC) (Goldreich, 1998), Federated Learning (Kairouz et al., 2021), or Homomorphic Encryption (Yi et al., 2014) can be leveraged to ensure that agents collaborate without compromising data privacy.
@@ 5. Conclusion and Future Work
In conclusion, this paper presents a roadmap for advancing LLM-based Multi-Agent systems in software engineering, outlining immediate and long-term research priorities. As part of this vision, we conducted a review of recent works in software engineering that apply LMA systems across various stages of the software development lifecycle. This review highlights key contributions in areas such as requirements engineering, code generation, software quality assurance, and maintenance. We further outline a structured research agenda to advance LMA integration in SE. Future work will concentrate on addressing the set of critical research questions identified, with a focus on enhancing LMA capabilities and optimizing their synergy within software development processes.
In the immediate term, efforts will be dedicated to enhancing the capabilities of LLM-based agents in representing specialized SE roles. This will involve creating specialized datasets and pre-training tasks that mirror the complex realities of SE tasks. Additionally, there will be a focus on formulating advanced prompting strategies, which can refine the agents’ cognitive functions and decision-making skills. Looking towards the longer-term objectives, the emphasis will transition towards optimizing the synergy between agents. The initial step involves examining optimal strategies for task allocation between humans and LLM-based agents, capitalizing on the unique strengths of both entities. Further, we need to delve into developing scalable methodologies for LMA systems, enabling LMA systems to efficiently orchestrate and complete large-scale, multifaceted SE projects. Furthermore, ensuring the privacy and confidentiality of data within LMA systems is critical. This entails exploring data management and access control mechanisms to protect sensitive information while still enabling the essential exchange of insights among project stakeholders. By systematically exploring these research questions, our goal is to drive innovation in LMA systems for software engineering, aiming to create a more cohesive, effective, and flexible LMA-driven development process.

### v3

@@ 5. Research Agenda
Previous research has laid the groundwork for the exploration of LMA systems in software engineering, yet this domain remains in its nascent stages, with many critical challenges awaiting resolution. In this section, we outline our perspective on these challenges and suggest research questions that could advance this burgeoning field. As illustrated in Figure 4, we envision two phases for the development of LMA systems in software engineering. We discuss each of these phases below and suggest a series of research questions that could form the basis of future research projects.
Figure 4. Research Agenda for LLM-Based Multi-Agent Systems in Software Engineering
@@ 5.1.1. Refining Role-Playing Capabilities in Software Engineering
The role-playing capabilities of LLM-based agents are pivotal within LMA systems (Wang et al., 2023d). To address the complexity of software engineering tasks, we need specialized agents capable of adopting diverse roles to tackle intricate challenges throughout the software development lifecycle.
Current State. Existing LMA systems, such as ChatDev (Qian et al., 2024c), MetaGPT (Hong et al., 2023), and AgileCoder (Nguyen et al., 2024), effectively simulate roles like generic software developers and product managers. The agents in these systems rely on general-purpose LLMs such as ChatGPT. Although LLMs like ChatGPT exhibit strong programming skills, they still lack the nuanced expertise required in SE (Hou et al., 2023). This limitation hampers their ability to simulate other SE-specific roles. For example, roles involving vulnerability detection or security auditing require a deep understanding of security protocols, threat modeling, and the latest vulnerabilities. However, multiple studies have identified deficiencies in ChatGPT’s ability to accurately detect and repair vulnerabilities (Fu et al., 2023; Sridhara et al., 2023; Chen et al., 2023b). This shortcoming underscores the need to integrate domain-specific expertise into LLMs to better support specialized software engineering roles.
Opportunities. To address this limitation, we propose a structured and actionable three-step approach encompassing the identification, assessment, and enhancement of role-playing abilities, which are:
@@
    Value Addition Modeling: The next crucial step is value addition modeling (Mendes et al., 2018), which evaluates the potential advantages that LLM-based agents could bring to each prioritized role. This process involves constructing detailed, data-driven models to analyze key performance indicators such as efficiency improvements, cost reductions, quality enhancements, and the acceleration of innovation resulting from the integration of agents. Pilot projects can be deployed to gather empirical data on these metrics when LLM-based agents are applied to specific tasks. Important factors to consider include the automation of repetitive tasks, the augmentation of human capabilities, and the inclusion of new functionalities that were previously unattainable. It is important to note that the value added by LLM-based agents can differ significantly across different domains; for example, roles in software development may prioritize automation, whereas domains like systems architecture might see more value in LLMs augmenting complex decision-making around resource allocation or performance optimization, where human expertise and contextual understanding remain essential. By quantifying these value propositions, organizations can allocate resources more strategically to roles where LLM-based agents are likely to yield the highest return on investment.
@@ 5.1.2. Advancing Prompts through Agent-oriented Programming Paradigms
Effective prompts are crucial for the performance of LLM-based agents. However, creating such prompts is challenging due to the need for a framework that is versatile, effective, and robust across diverse scenarios. Natural language, while flexible, often contains ambiguities and inconsistencies that LLMs may misinterpret. Natural language is inherently designed for human communication, where human communication relies on shared context and intuition that LLMs lack. In contrast, LLMs interpret text based on statistical patterns from large datasets, which may lead to different interpretations than those intended for humans (Sun et al., 2024; Zeng et al., 2022). This highlights the need for a specialized prompting language designed to augment the cognitive functions of LLM-based agents and treats LLMs as the primary audience. Such a language can minimize ambiguities and ensure clear instructions, resulting in more reliable and accurate outputs.
Current State. Multiple prompting frameworks are released to facilitate the usage of LLMs. For example, DSPy (Khattab et al., 2023) and Vieira (Li et al., 2024b) enable fully automated generation of prompts. AutoGen (Wu et al., 2023a) and LangChain (Mavroudis, 2024) support retrieval-augmented generation (RAG) (Gao et al., 2023) and agent-based workflows. However, these frameworks are still human-centered. They often prioritize human readability and developer convenience. As a result, there is a lack of research on a language that treats LLMs as the primary audience for prompts.
@@
Opportunities. Agent-oriented programming (AOP) (Shoham, 1993) offers a promising foundation for this approach. Similar to how Object-Oriented Programming (OOP) (Wegner, 1990) organize objects, AOP treats agents as fundamental units, focusing on their reasoning, objectives, and interactions. An AOP-based prompting language could enable the precise expression of complex tasks and constraints, allowing LLM-based agents to perform their roles with greater efficiency and accuracy. Extending this concept to Multi-Agent-Oriented Programming (MAOP) (Boissier et al., 2020; Bordini et al., 2009) allows for the creation of systems where multiple LLM-based agents can collaborate, communicate, and adapt to evolving contexts. By explicitly defining agent behaviors, communication patterns, and task hierarchies, we can reduce ambiguity, mitigate hallucinations, and improve task execution in LMA systems.
Further, such a prompting language must be expressive enough to handle diverse and complex tasks, yet simple enough for users to easily adopt. Conversely, overly simplified languages may lack the expressive power needed to represent complex software engineering workflows. A complex language that introduce a steep learning curve due to their syntax, hinder adoption, especially for users who require simpler interfaces for prompt creation and modification. Balancing functionality and usability will be another key research question to its success.
Additionally, this process may involve tailoring prompts specifically for different LLM models and their versions, as variations in model architectures, training data, and capabilities can affect how they interpret and respond to prompts. What works effectively for one model may not perform as well for another, necessitating careful adjustments. Current prompting languages lack mechanisms to easily adapt prompts across models, requiring manual adjustments and experimentation to achieve consistent performance.
While AOP-based prompting may not be the final solution, it represents an important step toward developing an AI-oriented language with grammar tailored specifically for LLMs. This new approach could further refine communication with LLM-based agents, reducing misinterpretation and significantly enhancing overall performance.
@@ 5.2.1. Human-Agent Collaboration
Optimally distributing tasks between humans and LMA agents to leverage their respective strengths is essential. Humans bring unparalleled creativity, critical thinking, ethical judgment, and domain-specific knowledge (Markauskaite et al., 2022). In contrast, LLM-based agents excel at rapidly processing large datasets, performing repetitive tasks with high accuracy, and detecting patterns that might elude human observers.
Current State. Several LMA systems incorporate human-in-the-loop designs. For instance, AISD (Zhang et al., 2024d) involves human input during requirement analysis and system validation, where users provide feedback on use cases, system designs, and prototypes. Similarly, MARE (Jin et al., 2024) leverages human assessment to refine generated requirements and specifications. Although these works demonstrate the feasibility of human contributions, key research questions, including optimizing human roles, enhancing feedback mechanisms, and identifying appropriate intervention points, are still underexplored.
Opportunity. Developing role-specific guidelines that outline when and how human intervention should occur is essential. These guidelines should assist in identifying critical decision points where human judgment is indispensable, such as ethical considerations, conflict resolution, ambiguity handling, and creative problem-solving. For example, ethical decisions necessitate human oversight to ensure alignment with societal norms and values, and conflict resolution may require negotiation skills that LLM-based agents lack.
To facilitate seamless collaboration, designing intuitive user-friendly interfaces and interaction protocols is essential (Wang et al., 2024a). Natural language interfaces and adaptive visualization techniques can make interactions more accessible. These interfaces should efficiently present agent outputs in a digestible format and collect user feedback, while also managing the cognitive load on human collaborators. It is important to note that these interfaces may need to be tailored differently for each human role, as the needs of a project manager, a software developer, and a quality assurance engineer will vary significantly.
Given the complexity of information generated during the agents’ workflows (Josifoski et al., 2023), designing such interfaces poses challenges. For instance, presenting modifications suggested by an agent at varying levels of abstraction ensures that each stakeholder can engage with the information at the right depth. A project manager might focus on the broader implications, such as the high-level impact on project timelines or deliverables, whereas a developer or architect might drill down into specific implementation details. Role-specific interfaces will be key to ensuring each stakeholder can effectively collaborate with the agents and extract the necessary information in a manner suited to their specific responsibilities.
@@ 5.2.2. Evaluating the LMA systems
Current State. Numerous complex benchmarks have been proposed to challenge the capabilities of LLMs in critical aspects of software development, such as code generation (Zhuo et al., 2024; Jain et al., 2024; Liu et al., 2024c). While these benchmarks have advanced the field by providing measurable metrics for individual tasks, their focus on isolated problem-solving reveals limitations as software engineering projects become more complex. Software engineering is inherently collaborative, with key activities like joint requirements gathering, code integration, and peer reviews playing essential roles in the process. Current benchmarks often overlook these aspects, failing to assess how well LLMs perform in tasks that require cooperation and collective decision-making.
Opportunities. There is a growing need for benchmarks that evaluate the cooperative abilities of LLMs in multi-agent settings, particularly for software engineering tasks. These benchmarks should simulate real-world collaborative scenarios where LLM agents work together to achieve common development goals.
@@
Current State. Existing LMA systems face significant challenges when scaling up to handle complex software projects. Our case studies illustrate these limitations clearly. For instance, ChatDev was unable to autonomously develop a functional Tetris game. In real-world projects with higher complexity, this limitation becomes even more pronounced.
Opportunities. First, as software projects grow in size and complexity, breaking down high-level requirements into manageable sub-tasks becomes more difficult. It is not just about handling more tasks but also managing the intricate interdependencies between them. A hierarchical task decomposition approach can help, where higher-level agents oversee broader objectives and delegate specific tasks to lower-level agents. This structure streamlines planning and makes global task allocation more efficient.
@@ 5.2.4. Leveraging Industry Principles
Current State. As we described in Section 3, numerous works (Qian et al., 2024c; Hong et al., 2023; Nguyen et al., 2024; Al-Saqqa et al., 2020) are designed using popular process models like the Waterfall and Agile. For example, ChatDev (Qian et al., 2024c) emulates a traditional Waterfall approach, breaking tasks into distinct phases (e.g., requirement analysis, design, implementation, testing), with agents dedicated to each phase. AgileCoder (Nguyen et al., 2024) incorporates the Agile methodologies, leveraging iterative development, continuous feedback loops, and collaborative sprints.
Opportunities. However, current LMA systems often do not leverage more specialized and modern industry practices, such as Value Stream Mapping, Design Thinking, or Model-Based Systems Engineering (MBSE). Additionally, frameworks like Domain-Driven Design (DDD), Behavior-Driven Development (BDD), and Team Topologies remain underutilized. These methodologies emphasize aligning development with business goals, improving user-centric design, and optimizing team structures—key components that could further enhance the efficiency, adaptability, and effectiveness of LMA systems.
@@
In the context of software development, predicting the optimal configuration for LMA systems at the outset is unrealistic due to the inherent complexity and variability of tasks (Leffingwell and Widrig, 2000). The dynamic nature of software requirements and the unpredictable challenges that arise during development necessitate systems that can adapt on the fly (Liu et al., 2023). For example, a sudden shift in project requirements or unexpected delays caused by dependencies on external components. Therefore, LMA systems must be capable of dynamically adjusting their scale, strategies, and structures throughout the development process.
Current State. Most existing LMA systems (Qian et al., 2024c; Hong et al., 2023) operate with static architectures characterized by fixed agent roles and predefined communication patterns. Recent research efforts (Liu et al., 2024e; Zhang et al., 2024g) have introduced mechanisms for adaptive agent team selection and task-specific collaboration strategies. These methods enable the selection of suitable agent team configurations for specific tasks, however, they still fall short of true dynamic adaptation and lack the capability to adjust to real-time changes. To the best of our knowledge, no previous work addresses the need for on-the-fly adjustments in response to evolving project demands.
Opportunities. To minimize redundant work, LMA systems should continuously evaluate existing solutions (Wang et al., 2024c), identifying reusable elements for new requirements. By learning from each development cycle, the system can recognize patterns of efficiency and inefficiency, enabling it to make informed decisions when handling similar tasks in the future or adapting existing solutions to new requirements.
@@
Current State. The challenge of ensuring privacy while managing partial information has been extensively studied in the field of computer security.  (Dinur and Nissim, 2003; Bennett et al., 1988). To the best of our knowledge, existing research has yet to provide a solution to these challenges for LMA systems in SE.
Opportunities. To address these challenges, robust and fine-grained access control mechanisms must be implemented across organizational boundaries. It is essential to prevent unauthorized access while still accommodating the varied data access needs of the system. Traditional models like Role-Based Access Control (RBAC) (Sandhu, 1998) and Attribute-Based Access Control (ABAC) (Hu et al., 2015) may need to be extended to handle the dynamic nature of multi-agent systems effectively. Establishing protocols that allow agents to share insights derived from sensitive data, without exposing the data itself, is critical. Advanced privacy-preserving techniques like Differential Privacy (Dwork, 2006), Secure Multi-Party Computation (SMPC) (Goldreich, 1998), Federated Learning (Kairouz et al., 2021), or Homomorphic Encryption (Yi et al., 2014) can be leveraged to ensure that agents collaborate without compromising data privacy.
@@
6. Discussion
6.1. A Comparison with the Mixture of Experts Paradigm
Another paradigm that has recently attracted much attention from both academia and industry is the Mixture of Experts (MoE) paradigm (Zhu et al., 2024; Cai et al., 2024). MoE organizes an LLM into multiple specialized components known as “experts.” Each expert is designed to focus on specialized tasks. Further, a gating mechanism is employed to dynamically activate the most relevant subset of experts based on the input. While MoE is promising, LMA systems offer several distinct advantages:
One limitation of MoE is its high resource consumption. MoE models contain multiple experts within a single architecture, which makes the total number of parameters rather huge. Furthermore, training MoE is more resource-intensive and time-consuming than standard LLMs. This is mainly due to the complex training process for the gating mechanism. Training the gating mechanism involves optimizing the selection process for the most relevant experts, which adds considerable overhead.
Since specific experts are dynamically activated based on input, MoE can be viewed as a method to learn the internal routing of LLMs. However, there is no interaction and communication between experts in MoE. On the other hand, LMA systems usually are designed to resemble real-world collaborative workflows. Agents in LMA systems can actively communicate with each other, exchange information, and iteratively refine the output based on feedback from other agents. More importantly, LMA systems can also integrate external feedback from tools such as compilers, static analyzers, or testing frameworks. LMA systems also facilitate seamless and continuous human-in-the-loop collaboration, enabling human experts to intervene, validate outputs, and provide guidance at any stage of the process. As a result, we consider LMA systems to be a more appropriate approach to MoE to address the multifaceted challenges of software engineering.
6.2. Threat to Validity
One potential threat to validity lies in the possibility of inadvertently excluding relevant studies during the literature search and selection process. To mitigate this risk, we conducted a comprehensive search on the DBLP database, ensuring coverage of a broad spectrum of studies, including preprints. Additionally, we enhanced the search process by combining automated querying with forward and backward snowballing, aiming to identify and include all pertinent studies.
@@ 7. Conclusion and Future Work
This paper explores the evolving role of LMA systems in shaping the future of Software Engineering 2.0 (Lo, 2023). To support this vision, we first present a systematic review of recent applications of LMA systems across different stages of the software development lifecycle. Our review highlights key advancements in areas such as requirements engineering, code generation, software quality assurance, and maintenance. To further understand the current landscape, we conduct two case studies that illustrate the practical uses and challenges of LMA systems. Based on these insights, we propose a structured research agenda aimed at advancing LMA integration in software engineering. Future work will focus on addressing critical research questions to enhance LMA capabilities and optimize their synergy with software development processes.
In the immediate term, efforts will be dedicated to enhancing the capabilities of LLM-based agents in representing specialized SE roles. This will involve creating specialized datasets and pre-training tasks that mirror the complex realities of SE tasks. Additionally, there will be a focus on formulating advanced prompting strategies, which can refine the agents’ cognitive functions and decision-making skills. Looking towards the longer-term objectives, the emphasis will transition towards optimizing the synergy between agents. The initial step involves examining optimal strategies for task allocation between humans and LLM-based agents, capitalizing on the unique strengths of both entities. Further, we need to develop scalable methodologies for LMA systems, which would enable them to orchestrate and complete large-scale, multifaceted SE projects efficiently. Moreover, ensuring the privacy and confidentiality of data within LMA systems is also critical. This entails exploring data management and access control mechanisms to protect sensitive information while still enabling the essential exchange of insights among project stakeholders. By systematically exploring these research questions, we aim to drive innovation in LMA systems for software engineering and create a more cohesive, effective, and flexible LMA-driven development process.